OpenStudy (anonymous):

what pattern do you discover when solving even numbered exponents with negative bases? A.solution is always positive B.solution is always negative C. solution depends on the sign of the exponent D.none of the above

OpenStudy (anonymous):

A

OpenStudy (mathstudent55):

Best thing to do is try a few and see if you can tell the pattern from the choices you have. (-1)^2 = 1 (-2)^2 = 4 (-5)^4 = 625 (-3)^(-2) = 1/9 (-1)^(-4) = 1